OpenStudy (anonymous):

Help with Algebra 2?

OpenStudy (anonymous):

Solve system by elimination.

OpenStudy (anonymous):

@IAmSinged

OpenStudy (anonymous):

x = -3, y = 6, z = -1 using a matrix solver I wrote. Check it using : http://www.bluebit.gr/matrix-calculator/linear_equations.aspx

OpenStudy (anonymous):

*x = 3, typo'd

OpenStudy (anonymous):

Im sorry I dont get it

OpenStudy (anonymous):

can you explain it a little more please?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@IAmSinged

OpenStudy (anonymous):

I solved the matrix using very high-level math, so I'll tell you a "dumber" way to solve it. The method is called Naive Gaussian Elimination, because if there are any numbers on the diagonal, it will fail completely. First, we turn the system into a matrix system Ax = b: A: -1 -1 -1 -4 4 5 2 0 2 B: -8 7 4 A is me taking everything before the equal sign and putting it into a matrix, and B is me taking everything right of the equal sign and putting it into a matrix. posting this now for you to have something to look at

OpenStudy (anonymous):

Continuing: So we take -1 from the first row of the matrix (aka our first equation) and what we want to do is to zero out the -4 and 2 below it. *NOTE: row and equation mean the same thing. How do we do this? We do this by adding row 1 multiplied by some constant to the second and third rows. What is this constant? This constant is whatever number multiplied by -1 will get me -4 in the case of 0'ing the 2nd row and whatever number multiplied by -1 gets us 2. For row 2 in our row matrix: We know that 4 * -1 = -4, so we know that our constant is 4. This tells us we are going to subtract row 1 from row 2 4 times. So we replace row 2 with: [-4 4 5] - 4[ -1 -1 -1] = [0 8 9] We do the same for our matrix B: 7 - (4 * -8) = 39 So now our A matrix is: -1 -1 -1 0 8 9 2 0 2 and our B matrix is: -8 39 4

OpenStudy (anonymous):

Continuing: For row 3: We know that -2 * -1 = 2, so we are going to subtract -1 * row 1 from row 3 twice, which basically means we're adding 2* row 1 to row 3. So for matrix A we do: row 3 = [ 2 0 2] + 2[ -1 -1 -1] = [0 -2 0] likewise, we do the same for matrix B: row 3 = 4 - 16 = -12 So now our matrix A is: -1 -1 -1 0 8 9 0 -2 0 And our B matrix is: -8 39 -2 This is equivalent to saying that our system of equations is now: -x-y-z = -8 8y + 9z = 39 -2y = -12 We ignore the 3rd element in the diagonal, because under the last element of a diagonal in a square matrix, there is nothing below it to 0 out. Note for that specific case we have an equation with 1 variable, which is -2y = -12. This gets us y = 6. Substituting y =6 into the second equation which has 2 variables, we get: 48 + 9z = 39 9z = -9 z = -1 So now we know that y = 6 and z = -1. Plugging into the first equation, we get x = 3. So now we have all 3 variables: x = 3, y =6, z = -1 This system of equations is kind of weird for NGE, since we have 0 on the diagonal, which normally means NGE fails. However, we ended up with an equation that only has y, so it ended up working out. A video is probably much better than a wall of text if you mind just shuts down from looking at walls of text: NGE in general: https://www.youtube.com/watch?v=m3ooPCNaoTA